if k is an integer then which of the following is true??????????/


a)an integer of the form 4k+3 can always be put in the form of 2k+1


b) an integer of the form 2k-1 can always be put in the form of 4k+1


c) an integer of the form 2k-1 can always be put in the form of 4k+3


d) an integer of the form 2k+1 can be always put in the form 4k+3


e) an integer of the form 4k+1 can always be put in the form 2k-1


 


sir plzzz explain each and every step...........    i will surely approve..................... need the solution urgently...............


 


kedar joshi....

Grade:Upto college level

2 Answers

Shivam Bhagat
12 years ago

a b and e are true

Nimon B
12 years ago

a) 4k+3 can be written as 4k+2+1=2(2k+1)+1 = 2c+1 where c=2k+1;i.e, some canstant. Correct

b) 2k-2+1=2(k-1)+1=2c+1 which is not equal to 4k+1 since multiplying 2c+1 by 2, we get 4c+2. Incorrect

c) 2k-1=2k-2+1=2(k-1)+1=2c+1 and by a), we get 2c+1=4k+3. Correct

d) reverse of part a). Correct

e) reverse of part b). Incorrect